In a sweepstakes sponsored by Gemini Paper Products, 100,000 entries have been received. If 2 grand prize, 10 first prizes, 35 second prizes, and 600 third prizes are to be awarded, what is the probability that a person who has submitted one entry will win a prize?
A) The probability is 0.00717
B) The probability is 0.01347
C) The probability is 0.00647
D) The probability is 0.00654

Answer by stanbon(75887)

In a sweepstakes sponsored by Gemini Paper Products, 100,000 entries have been received. If 2 grand prize, 10 first prizes, 35 second prizes, and 600 third prizes are to be awarded, what is the probability that a person who has submitted one entry will win a prize?
---------------------
# of prizes: 2+10+35+600 = 647
---
P(win a prize) = 647/100,000 = 0.00647
==========================================
Cheers,
Stan H.
==================================
A) The probability is 0.00717
B) The probability is 0.01347
C) The probability is 0.00647
D) The probability is 0.00654